OpenStudy (anonymous):

How do I calculate the sum of the infinite series: (4/9)^3 - (4/9)^4 + (4/9)^5 - (4/9)^6 + ... Giving the answer as a fraction?

OpenStudy (anonymous):

Chris do you know some LaTeX reader ? I wanted to practice LaTeX. I tried MikTex but doesn't work on my PC.

OpenStudy (anonymous):

For the question anaria take - terms and plus terms then sum (G.P sum) and then normal subtraction

OpenStudy (anonymous):

sum of an infinite geometric series is: S=a/(1-r) = (4/9)^3/(1-4/9) =64/405 =

OpenStudy (anonymous):

raheen ... thanks. looks right to me. :)

OpenStudy (dumbcow):

geometric series a1 = (4/9)^3 r = -4/9 \[s = \frac{a_{1}}{1-r} = \frac{(4/9)^{3}}{1+4/9} = \frac{4^{3}}{9^{3}}*\frac{9}{13} = \frac{64}{81*13} \approx 0.06\]

OpenStudy (anonymous):

ishaan that would be ok if not an infinite sequence

OpenStudy (anonymous):

wish that anaria

OpenStudy (anonymous):

yeah your right chris and raheen did that correct

OpenStudy (dumbcow):

r should be negative, making denominator (1+4/9) making the sum less

OpenStudy (anonymous):

Ishaan ... I am not Chris. I think you must have me confused with someone else. Confusingly, dumbcow also has an option 64/1053 which is also one of the options! Now which one is right. lol

OpenStudy (anonymous):

chris (dumbcow).........he did that right

OpenStudy (dumbcow):

If a series in alternating sign, then the common ratio must be negative :)

OpenStudy (anonymous):

Ok guys thanks ... :)